# Retrieving Semantics from the Deep: an RAG Solution for Gesture Synthesis

### v1

@@ 1 Introduction
McNeill [44, 43] categorizes gestures intoSemantic gestures are context-driven and bear a specific meaning that complements the utterance [43, 6]. Rule-based approaches retrieve these semantic gestures and combine them for generation, while neural methods leverage the data to learn the synthesis process. However, the former tends to result in an unnatural outcome and the latter struggles to generate semantic gestures because these gestures, while being a part of the existing datasets, occur rather sporadically compared to the rhythmic beat gestures [52].
@@
To solve the specification problem, we explicitly retrieve the relevant exemplar motions from a gesture database using either of the two retrieval algorithms proposed in Sec. 3.4. More specifically, we show that the chain-of-thought reasoning abilities of Large Language Models [53, 73] can be exploited to extract which parts of the utterance would likely be semantically gestured, and what ‘type’ of semantic gesture (iconic, deictic, metaphoric, etc) the phrase would invoke. We also demonstrate that one could, alternatively, use specific linguistic elements to base the retrieval on. In particular, we focus on discourse connectives such as because, while, and on one hand, which have been shown to affect the gestural patterns [45, 9].
The animation problem, in our context, requires integrating the retrieved gesture motions (corresponding to a subset of the sentence) into the overall motion generated for the full sentence/utterance (see Fig. 1). To this end, we design an inference-time approach to augment the gesture generation capability of diffusion models with retrieved gestures clips. Our approach allows us to surgically integrate the semantic gestures into the generation process of the base model, while not affecting the synthesis for the remaining parts of the utterance.
@@ 2.1 Co-Speech Gesture Synthesis
Co-speech gestures consist of hand and arm movements that occur in sync with speech and help convey meaning and structure discourse. Earlier methods in co-speech gesture generation [10, 61] perform well on semantics and appropriateness of gestures [52]. These systems are intent-driven and utilize heuristics to retrieve the most appropriate gesticulation units for given speech, resulting in high semantic alignment. However, since they insert predefined units into animation, the synthesis quality of these systems tends to be unnatural. To mitigate the lack of naturalness, recent learning-based methods [34, 2, 18, 24] employ deep networks to convert speech input to gesture output and aim to ensure smooth motions. These methods are mostly data-driven and rely on learning speech-to-gesture matching through large scale gesture datasets [39, 65, 17, 20]. Most of these recent methods [66, 34, 1, 64] suffer from reduced communicative efficacy and low semantic alignment, due to bad generalization to semantic co-speech gestures in the dataset. Furthermore, generation and reproduction of semantically appropriate gestures corresponding to speech still remains a challenge for these systems [67].
A few recent approaches directly addressed this problem. Kucherenko et al. [35] use speech to predict the gesture properties, which inform a probabilistic model on what type of gesture to generate. In comparison, our framework proposes to use RAG to follow the gesture type information instead of taking gesture properties as conditioning input. Liang et al. [38] decouple semantic irrelevant cues like rhythmic beats and volume to generate plausible gestures. Mughal et al. [47] generate semantic gestures by providing word-level semantic control and demonstrate how their diffusion model uses it to sample more semantically meaningful gestures. SemanticGesticulator [71] is a generative retrieval framework which retrieves semantic gestures from a curated motion library and aligns them with rhythmic gestures. Compared to their approach, ours does not rely on a curated library. Rather, it introduces linguistics-based explicit algorithms to retrieve semantics from the existing data [39]. Moreover, we perform retrieval augmentation using a diffusion model at inference time rather than fine-tuning our gesture generator to follow the retrieval.
@@ 2.2 Retrieval Paradigms for Generative Methods
Employing database retrieval for improved animation performance is a prevalent paradigm in motion and gesture synthesis. Classical [8, 11] and learning-based motion matching approaches [28, 29] have been applied for better character control in motion synthesis [58]. Diffusion-based text-to-motion synthesis frameworks have also dealt with the problem of semantic control at a global level [68] or at the local keyframe level [23, 30]. Zhang et al. [68] introduce RAG to diffusion-based motion synthesis by using retrieval based on global text similarity for text-to-motion tasks. In contrast to the text-to-motion task, gesture generation is affected by language content and prosody at varying levels of semantic granularity. That means one can neither perform global semantic matching between the retrieved and generated motion [68] nor simply copy and paste retrieved motion parts onto the keyframe locations [23] without considering the differences in speech contexts.
@@ 2.3 The effect of Linguistic Dynamics on Gestures
The goal of our retrieval approach is to extract semantic gestures from the database. Our proposed design is therefore motivated by the relation between language dynamics and co-speech gestures. Laparle et al. [36] demonstrate how gesture features like hand shape or orientation are important to distinguish between topics in discourse. Gestures can also indicate the relations between different parts of the message. For example, a contrast relation can be expressed by the lexical marker on the one hand…on the other hand, framing the two contrasting arguments as items placed on separate hands [26]. The raised index finger draws the attention of the interlocutor to new and important topics, and is commonly used to express exceptions or concessions (which can be expressed linguistically using however) [7, 32, 31].
@@ 3.1 Gesture Generation with Latent-Diffusion
As the first step, we build a robust generation framework based on latent-diffusion modeling [55], which serves as our base model to generate plausible co-speech gesture sequences. The gesture sequence $\mathbf{x}\in\mathbb{R}^{N\times d_{\mathbf{x}}}$ consists of $N$ frames of human motion with $d_{\mathbf{x}}$ representing a combination of $J$ full body joints represented using the 6D rotation representation of [74], FLAME model parameters for face motion, root translation and foot contact labels [40]. Instead of training our diffusion framework on raw motion $\mathbf{x}$, we encode motion using a VAE encoder, which provides separate encodings for different body parts [47]. These part-wise VAE encodings are then used to train our conditional latent-diffusion model that generates co-speech gestures.
@@ 3.2 Latent Initialization through DDIM Inversion
So far, our base neural generation framework is purely data-driven and lies on one end of neural-to-explicit spectrum. It resembles the majority of state-of-the-art gesture generation frameworks [72, 51, 3], which rely on learned patterns to sample relevant gestures for corresponding speech and struggle to sample the semantic ones. On the other hand of the spectrum, pure retrieval based approaches can naïvely “paste” raw exemplar motions onto the generated ones which hurts the naturalness of motion. For instance, one could use diffusion inpainting [41] to force our diffusion framework to follow the retrieval (See Sec. 4.3 for analysis), achieving sub-par results. Therefore, we carefully infuse the retrieved gestures into the diffusion sampling process by performing the retrieval transfer in the latent space of the trained diffusion model and provide better sampling path for its generation through retrieval guidance (Sec. 3.3). This preserves the quality of the base model, adapts it to generate semantically rich gestures and provides further control over the influence of retrieval insertion.
In our framework, all retrieval algorithms (discussed later in Sec. 3.4) must output an exemplar motion $\mathbf{m}$, which can be encoded as the retrieval $\mathbf{r}\in\mathbb{R}^{M\times d}$ through our gesture encoder (Sec. 3.1). We also have access to the text and audio corresponding the to retrieved motion. Our goal is to extract the motion chunk for the retrieved semantic gesture around a word of interest (e.g. ‘because’ in Fig. 3) and insert it into the motion around the corresponding text in the input text. Specifically, let the retrieved motion contain the semantically relevant gesture chunk in the frame window $(s_{\text{retr}},e_{\text{retr}})$ for the marked word. The gesture in this window must be transferred onto the corresponding word location $(s_{\text{query}},e_{\text{query}})$ in the to-be-generated sequence, also referred to as query $\mathbf{\hat{z}}$ (see Fig. 3).
@@ 3.3 Retrieval Guidance
To control the amount of guidance at each step, one can control the number of iterations for latent update at each diffusion timestep $t$. As an example, the generation from $\mathbf{\hat{z}}_{\text{retr}}^{(0)}$ without any guidance will be the weakest form of retrieval augmentation where gestures are constrained to match only at the starting timestep $t=0$. On the other hand, if one were to make sure retrieval is followed to its strongest extent, $\mathbf{\tilde{z}}_{\text{retr}}^{(t)}$ can be updated at each diffusion timestep $t$ such that $G_{retrieval}\rightarrow 0$. Similarly, one can perform more latent updates when $t$ is closer to $T$ and lesser or no updates when $t\rightarrow 0$, in order to only perform guidance at the start of the sampling process. Refer to Sec. 4.3 for analysis.
@@ 3.4 Retrieval for Semantic Gesture Extraction
Having discussed how to insert a retrieved gesture chunk into the gesture generation process, we now discuss retrieval algorithms that utilize linguistic context to determine which words in the query utterance are likely to invoke semantic gestures and what are those gestures. These algorithms retrieve semantically relevant gestures using domain knowledge of language and gesture type information However, our approach is not limited to these algorithms and it can be extended using any retrieval paradigm for gestures.
At the core, each retrieval algorithm ranks examples in a database (the BEAT2 dataset [40] in our case) according to a specific focus. During ranking, some of the steps are common across algorithms and are driven by over-arching themes in gesture understanding. For example, examples from the same speaker typically have less speaker-dependent variation and should be ranked higher. Another common step across algorithms is ranking based on prosodic prominence extraction [16]. Certain elements of speech typically stand out due to variations in pitch, loudness, or duration, referred to as prosodic prominence  [62, 60]. Prosodic prominence and gestures are closely connected: co-speech gestures often align with prosodically stressed parts of speech [54]. Therefore, we identify utterances in the database that match the prominence values of the marked word in the input speech [59]. An overview of individual steps is provided in Fig. 4(a).
@@
One of the commonly used classification for semantic gestures divides these gestures into iconic, metaphoric and deictic gestures [43]. Kucherenko et al. [35] demonstrate the benefit of utilizing these gesture properties in the synthesis framework, as gesture type information can decouple the specification problem from the animation. Therefore, we leverage the reasoning abilities of an LLM [53] and prompt it to identify words which may elicit a semantically meaningful gesture. We also use it to predict the gesture type on each predicted word. This results in a word-to-gesture type mapping, which we can use to retrieve examples from the database that contains type labels. Details on prompts and LLM are given in the supplemental material.
@@
A text has a structure that organizes its information into a coherent flow. This structure is achieved through discourse relations, semantic-pragmatic links such as cause-consequence and contrast, which hold between clauses and sentences [42, 56]. These relations can be signaled by lexical cues referred to as connectives, such as on the other hand or because. For example, in an utterance from BEAT2 dataset [40]: “I’ll go shopping if I’m not that tired”, the word “if” marks the conditional relation between two sentence segments. Prior works have demonstrated that discourse connectives have an effect on the gestural patterns [45, 9].
@@ 1 Speaker All Speakers
Table 1: Comparison with state-of-the-art methods trained on BEAT. We demonstrate superior performance, especially when generalizing well across multiple speaker identities.
@@ 4 Experiments
We first compare our approach with neural state-of-the-art baselines through quantitative (Sec. 4.1) and perceptual evaluation (Sec. 4.2). Further, we demonstrate how our RAG approach improves upon the diffusion-based RAG baseline. Lastly, we validate our method design through extensive ablations and user studies (Sec. 4.3).
@@ 4.1 Quantitative Evaluation
We evaluate our method on established metrics: Beat-Alignment [37], FID [65], L1 Divergence and Diversity, which measure different aspects of the motion quality. We tabulate the quantitative results in Tab. 1, where the results are divided on the basis of number of speakers used for training and evaluation. We observe that our approach achieves state-of-the-art metric performance when trained with all speakers. This also entails that our RAG-based diffusion approach generalizes and scales well across large-scale data with multiple speaker identities. For the case of single speaker training, we observe best performance in terms of diversity and L1 Divergence and second-best in Beat Alignment score. However, our diffusion-based method has lower FID compared to transformer and LSTM based methods [40, 39] that utilize ground-truth seed and can be over-fitted to a small-scale data of single speaker, while not being able to generalize to all speakers.
@@ 4.2 Perceptual Evaluation
Evaluating gesture synthesis performance is a highly challenging task due to the stochasticity in gesture production and subjective nature of gesture perception [63, 52]. Therefore, we perform three different user studies for comparison and ablative analysis (Sec. 4.3). First, we evaluate the synthesis quality compared to the baseline methods. We investigate the quality of gestures based on well-established measures of naturalness and appropriateness in relation to speech [48]. We ask the participants to perform pairwise comparisons between generations from baselines and our method. Results compare RAG-Gesture with discourse and LLM-based retrievals against state-of-the-art methods (Fig. 5). We observe that our method achieves better perceptual quality for both naturalness and appropriateness. For LLM based approach, we are marginally below the ground-truth preference, underscoring that our method produces not only natural, but also highly appropriate gestures.
@@ 4.3 Ablative Analysis
One of our RAG-driven baselines (ReMoDiffuse [68]) follows global text similarity for retrieval and proposes to train the diffusion network to follow the retrievals. In contrast, we design our retrieval to match local context around the semantically important word for gesticulation and do not require training/fine-tuning for RAG unlike existing approaches [71, 68]. We evaluate these differences by comparing our RAG-Gesture (Discourse) with (i) base ReMoDiffuse as well as (ii) a modified version of their diffusion architecture, which uses gesture encodings as input instead of raw motion, along with the proposed discourse-based retrieval algorithm for RAG (see supplementary for details). Note that this architecture is similar to SemanticGesticulator [71] at a higher level, which also performs retrieval merging in encoded gestures during fine-tuning.
@@
Firstly, we observe that our modified version of ReMoDiffuse which uses local semantic retrieval during training performs better than the base model, both in terms of perceptual quality and metrics. Moreover, we also see higher faithfulness to the semantic retrievals through low MPJPE which entails that properly specifying the retrieved motion on the exact word affects the gesture quality. Secondly, our inference-time RAG-Gesture is better than the training-based model, even after utilizing our retrieval algorithm. This underscores our method’s superiority and the importance of having a good gesture animation performance.
@@
For evaluating perceptual differences between LLM-based gesture type retrieval and discourse-based retrieval, we conduct an additional user study where we ask the participants to evaluate differences between generated gestures from two algorithms in terms of naturalness and appropriateness (Fig. 7).
Figure 7: Study Preference comparing LLM and Discourse algorithms.
@@
Since our approach can control the influence of the retrieved gesture onto the generated one, we analyze how much influence is optimal to synthesize semantically accurate yet natural looking gestures. We compare RAG-Gesture’s gesture prediction against varying degrees of retrieval insertion (Tab. 3), where “Inpainting” means naïvely pasting the retrieval over the gesture during synthesis and “No RAG” represents no retrieval insertion. We conduct another user study with a third “Retrieval Appropriateness” question which tries to gauge how semantically appropriate the generated gesture was during the insertion window Fig. 8.
@@
No RAG $0.519$ $0.447$ $8.644$ $59.6$
RAG by Inpainting [41] $0.446$ $0.450$ $8.457$ $28.1$
RAG by LI only $0.589$ $0.486$ $9.228$ $54.6$
RAG-Gesture (LI+RG) $0.447$ $0.471$ $9.038$ $35.0$
@@
Interestingly, we observe increasing preference of our retrieval-guided model as we move towards both extremes of the insertion. In terms of gesture perception, evaluators neither prefer generation without RAG (purely data-driven) nor completely strong insertion through Inpainting (purely retrieval-based). Lastly, our model achieves good balance of retrieval following in terms of MPJPE, prosodic alignment and diversity.

### v2

@@ 1 Introduction
McNeill [44, 43] categorizes gestures into beat gestures, which are rhythmic movements driven by prosody, and semantics-driven gestures, which carry the communicative intent. Semantic gestures are context-driven and bear a specific meaning that complements the utterance [43, 6]. Rule-based approaches retrieve these semantic gestures and combine them for generation, while neural methods leverage the data to learn the synthesis process. However, the former tends to result in an unnatural outcome and the latter struggles to generate semantic gestures because these gestures, while being a part of the existing datasets, occur rather sporadically compared to the rhythmic beat gestures [52].
@@
To solve the specification problem, we explicitly retrieve the relevant exemplar motions from a gesture database using either of the two retrieval algorithms proposed in Sec. 3.4. More specifically, we show that the chain-of-thought reasoning abilities of Large Language Models [53, 73] can be exploited to extract which parts of the utterance would likely be semantically gestured, and what ‘type’ of semantic gesture (iconic, deictic, metaphoric, etc.) the phrase would invoke. We also demonstrate that one could, alternatively, use specific linguistic elements to base the retrieval on. In particular, we focus on discourse connectives such as because, while, and on one hand, which have been shown to affect the gestural patterns [45, 9].
The animation problem, in our context, requires integrating the retrieved gesture motions (corresponding to a subset of the sentence) into the overall motion generated for the full sentence/utterance (see Fig. 1). To this end, we design an inference-time approach to augment the gesture generation capability of diffusion models with retrieved gesture clips. Our approach allows us to surgically integrate the semantic gestures into the generation process of the base model, while not affecting the synthesis for the remaining parts of the utterance.
@@ 2.1 Co-Speech Gesture Synthesis
Co-speech gestures consist of hand and arm movements that occur in sync with speech and help convey meaning and structure discourse. Earlier methods in co-speech gesture generation [10, 61] perform well on semantics and appropriateness of gestures [52]. These systems retrieve the most appropriate gesticulation units for given speech, resulting in high semantic alignment. However, since they insert predefined units into animation, the synthesis quality of these systems tends to be unnatural. To mitigate the lack of naturalness, recent learning-based methods [35, 2, 18, 24] employ deep networks to convert speech input to gesture output and aim to ensure smooth motions. These methods are mostly data-driven and rely on learning speech-to-gesture matching through large scale gesture datasets [39, 65, 17, 20]. Most of these recent methods [66, 35, 1, 64] suffer from reduced communicative efficacy and low semantic alignment, due to bad generalization to semantic co-speech gestures in the dataset. Furthermore, generation and reproduction of semantically appropriate gestures corresponding to speech still remains a challenge for these systems [67].
A few recent approaches directly addressed this problem. Kucherenko et al. [36] use speech to predict the gesture properties, which inform the model on what type of gesture to generate. Instead, our framework proposes to use RAG to follow the gesture type information instead of taking the properties as conditioning input. Mughal et al. [47] generate semantic gestures by providing word-level semantic control and use it to sample semantically meaningful gestures. Zhi et al. [74] initialize diffusion-based generation with text-aligned embeddings by adding noise to them (similar to inpainting [41]). In contrast, our approach inverts semantic exemplars into the diffusion latent space and uses them to initiate the denoising process. SemanticGesticulator [71] is a generative framework which retrieves semantic gestures from a curated motion library and aligns them with rhythmic gestures through a training-based approach. In comparison, our framework performs inference-time RAG using a diffusion model rather than fine-tuning the gesture generator to follow the retrieval. Secondly, our approach also utilizes prosody information for every retrieval instead of only using textual transcription. Thirdly, instead of relying on a curated library, our approach introduces explicit algorithms to retrieve semantics from the existing data [39]. Lastly, our framework is flexible enough to leverage any retrieval algorithm at inference time.
@@ 2.2 Retrieval Paradigms for Generative Methods
Employing database retrieval for improved animation performance is a prevalent paradigm in motion and gesture synthesis. Classical [8, 11] and learning-based motion matching approaches [28, 29] have been applied for better character control in motion synthesis [58]. Diffusion-based text-to-motion synthesis frameworks also deal with the problem of semantic control at a global level [68] or at the local keyframe level [23, 30]. Zhang et al. [68] apply RAG to diffusion-based text-to-motion synthesis using global text similarity for retrieval. Unlike text-to-motion, gesture generation is affected by language content and prosody at varying levels of semantic granularity. That means one can neither perform global semantic matching between the retrieved and generated motion [68] nor simply copy and paste retrieved motion parts onto the keyframe locations [23] without considering the differences in context.
@@ 2.3 The effect of Linguistic Dynamics on Gestures
The goal of our retrieval approach is to extract semantic gestures from the database. Our proposed design is therefore motivated by the relation between language dynamics and co-speech gestures. Laparle [37] demonstrates how gesture features like hand shape or orientation are important to distinguish between topics in discourse. Gestures can also indicate the relations between different parts of the message. For example, a contrast relation can be expressed by the lexical marker on the one hand…on the other hand, framing the two contrasting arguments as items placed on separate hands [26]. The raised index finger draws the attention of the interlocutor to new and important topics, and is commonly used to express exceptions or concessions (which can be expressed linguistically using however) [7, 32, 31].
@@ 3.1 Gesture Generation with Latent-Diffusion
As the first step, we build a robust generation framework based on latent-diffusion modeling [55], which serves as our base model to generate plausible co-speech gesture sequences. The gesture sequence $\mathbf{x}\in\mathbb{R}^{N\times d_{\mathbf{x}}}$ consists of $N$ frames of human motion with $d_{\mathbf{x}}$ representing a combination of $J$ full body joints represented using the 6D rotation representation of [75], FLAME model parameters for face motion, root translation and foot contact labels [40]. Instead of training our diffusion framework on raw motion $\mathbf{x}$, we encode motion using VAE encoders, which provide separate encodings for different body parts [47]. These part-wise VAE encodings are then used to train our conditional latent-diffusion model that generates co-speech gestures.
@@ 3.2 Latent Initialization through DDIM Inversion
So far, our base neural generation framework is purely data-driven and lies on one end of neural-to-explicit spectrum. It resembles the majority of state-of-the-art gesture generation frameworks [72, 51, 3], which rely on learned patterns to sample relevant gestures for corresponding speech and struggle to sample the semantic ones. On the other hand of the spectrum, pure retrieval based approaches can naïvely “paste” raw exemplar motions during the generation process which hurts the naturalness of motion. For instance, one could simply add noise to the retrieval and generate the rest through diffusion-based sampling [41], which forces the generation to exactly follow the retrieval, achieving sub-par results (see Sec. 4.3 for analysis). Therefore, we carefully infuse the retrieved gestures into the inference process by performing the retrieval transfer in the latent space of the trained diffusion model and provide a better sampling path for its generation through retrieval guidance (Sec. 3.3). This preserves the quality of the base model, adapts it to generate semantically rich gestures and provides further control over the influence of retrieval insertion.
In our framework, all retrieval algorithms (discussed later in Sec. 3.4) must output an exemplar motion $\mathbf{m}$, which can be encoded as the retrieval $\mathbf{r}\in\mathbb{R}^{M\times d}$ through our gesture encoder (Sec. 3.1). We also have access to the text and audio corresponding to the retrieved motion. Our goal is to extract the motion chunk for the retrieved semantic gesture around a word of interest (e.g. ‘because’ in Fig. 3) and insert it into the motion around the corresponding words in the input text. Specifically, let the retrieved motion contain the semantically relevant gesture chunk in the frame window $(s_{\text{retr}},e_{\text{retr}})$ for the marked word. The gesture in this window must be transferred onto the corresponding word location $(s_{\text{query}},e_{\text{query}})$ in the to-be-generated sequence, also referred to as query $\mathbf{\hat{z}}$ (see Fig. 3).
@@ 3.3 Retrieval Guidance
To control the amount of guidance at each step, one can control the number of iterations for latent update at each diffusion timestep $t$. As an example, the generation from $\mathbf{\hat{z}}_{\text{retr}}^{(T)}$ without any guidance will be the weakest form of retrieval augmentation where gestures are constrained to match only at the starting timestep $t=T$. On the other hand, if one were to make sure retrieval is followed to its strongest extent, $\mathbf{\tilde{z}}_{\text{retr}}^{(t)}$ can be updated at each diffusion timestep $t$ such that $G_{retrieval}\rightarrow 0$. Similarly, one can perform more latent updates when $t$ is closer to $T$ and lesser or no updates when $t\rightarrow 0$, in order to only perform guidance at the start of the sampling process. Refer to Sec. 4.3 for analysis.
@@ 3.4 Retrieval for Semantic Gesture Extraction
Having discussed how to insert a retrieved gesture chunk into the gesture generation process, we now discuss retrieval algorithms that utilize linguistic context to determine which words in the query utterance are likely to invoke semantic gestures and what are those gestures. These algorithms retrieve semantically relevant gestures using domain knowledge of language and gesture type information. However, our approach is not limited to these algorithms and it can be extended using any retrieval paradigm for gestures.
At the core, each retrieval algorithm ranks examples in a database (the BEAT2 dataset [40] in our case) according to a specific focus. During ranking, some of the steps are common across algorithms and are driven by over-arching themes in gesture understanding. For example, motion exemplars from the same speaker typically have less speaker-dependent variation and should be ranked higher. Another common step across algorithms is ranking based on prosodic prominence extraction [16]. Certain elements of speech typically stand out due to variations in pitch, loudness, or duration, referred to as prosodic prominence  [62, 60]. Prosodic prominence and gestures are closely connected: co-speech gestures often align with prosodically stressed parts of speech [54]. Therefore, we identify utterances in the database that match the prominence values of the marked word in the input speech [59]. An overview of individual steps is provided in Fig. 4(a).
@@
One of the commonly used classification for semantic gestures divides them into iconic, metaphoric and deictic gestures [43]. Kucherenko et al. [36] demonstrate the benefit of utilizing these gesture properties in the synthesis framework, as gesture type information can decouple the specification problem from the animation. Therefore, we leverage the reasoning abilities of an LLM [53] and prompt it to identify words which may elicit a semantically meaningful gesture. We also use it to predict the gesture type on each predicted word. This results in a word-to-gesture type mapping, which we can use to retrieve examples from the database that contains type labels. Details on prompts and LLM are given in the supplemental material.
@@
A text has a structure that organizes its information into a coherent flow. This structure is achieved through discourse relations, semantic-pragmatic links such as cause-consequence and contrast, which hold between clauses and sentences [42, 56]. These relations can be signaled by lexical cues referred to as connectives, such as on the other hand or because. For example, in an utterance from BEAT2 dataset [40]: “I’ll go shopping if I’m not that tired", the word “if" marks the conditional relation between two sentence segments. Prior works have demonstrated that discourse connectives have an effect on the gestural patterns [45, 9].
@@ 1 Speaker All Speakers
Table 1: Comparison with state-of-the-art methods trained on BEAT2. We demonstrate superior performance, especially when generalizing across multiple speaker identities.
@@ 4 Experiments
We first compare our approach with neural state-of-the-art baselines through quantitative (Sec. 4.1) and perceptual evaluation (Sec. 4.2). Further, we demonstrate how our RAG approach improves upon the training-based RAG baselines. Lastly, we validate our method design through extensive ablations and user studies (Sec. 4.3).
@@ 4.1 Quantitative Evaluation
We evaluate our method on established metrics: Beat-Alignment [38], FID [65], L1 Divergence and Diversity, which measure different aspects of the motion quality. We tabulate the quantitative results in Tab. 1, where the results are divided on the basis of number of speakers used for training and evaluation. We observe that our approach achieves state-of-the-art metric performance when trained with all speakers. This also entails that our RAG-based diffusion approach generalizes and scales well across large-scale data with multiple speaker identities. For the case of single speaker training, we observe best performance in terms of diversity and L1 Divergence and second-best in Beat Alignment score. However, our diffusion-based method has higher FID compared to transformer and LSTM based methods [40, 39] that utilize ground-truth seed and can be over-fitted to a small-scale data of single speaker, while not being able to generalize to all speakers.
@@ 4.2 Perceptual Evaluation
Evaluating gesture synthesis performance is a highly challenging task due to the stochasticity in gesture production and subjective nature of gesture perception [63, 52]. Therefore, we perform four different user studies for comparison and ablative analysis (Sec. 4.3). First, we evaluate the synthesis quality compared to the baseline methods. We investigate the quality of gestures based on well-established measures of naturalness and appropriateness in relation to speech [48]. We ask the participants to perform pairwise comparisons between generations from baselines and our method. Results compare RAG-Gesture with discourse and LLM-based retrievals against state-of-the-art methods (Fig. 5). We observe that our method achieves better perceptual quality for both naturalness and appropriateness. For LLM based approach, we are marginally below the ground-truth preference, underscoring that our method produces not only natural, but also highly appropriate gestures.
@@ 4.3 Ablative Analysis
ReMoDiffuse [68] uses global text similarity for retrieval and trains the diffusion network to follow it. In contrast, we design our retrieval to match local context around the semantically important word and do not require training/fine-tuning for RAG unlike existing approaches [71, 68].
Besides comparing these differences with ReMoDiffuse, we also evaluate RAG-Gesture (Discourse) against a training-based approach where we incorporate our discourse-based retrievals with ReMoDiffuse’s generator through the local retrieval merging strategy of SemanticGesticulator [71] (see supplementary for details). Note that this architecture is trained with encoded gestures instead of raw motion and is similar to SemanticGesticulator [71] at a higher level, which also performs retrieval merging with encoded gestures during fine-tuning.
@@
Firstly, we observe that the training-based RAG which uses local semantic retrieval during training performs better than ReMoDiffuse, both in terms of perceptual quality and metrics. Moreover, we also see higher faithfulness to the retrievals through low MPJPE which entails that properly specifying the retrieved motion on the exact word affects the gesture quality. Secondly, our inference-time RAG-Gesture is better than the training-based model, even after utilizing our retrieval algorithm. This underscores our method’s superiority and the importance of having a good gesture animation performance.
@@
For evaluating perceptual differences between LLM-based gesture type retrieval and discourse-based retrieval, we conduct an additional user study where we ask the participants to evaluate differences between generated gestures from the two algorithms in terms of naturalness and appropriateness (Fig. 7).
Figure 7: User Study comparing LLM and Discourse algorithms.
@@
Since our approach can control the influence of the retrieved gesture onto the generated one, we analyze how much influence is optimal to synthesize semantically accurate yet natural looking gestures. We compare RAG-Gesture’s prediction against varying degrees of retrieval insertion (Tab. 3) using Discourse-based retrieval. “Inpainting” means adding noise to the retrieval and synthesizing the remaining motion and “No RAG” represents no retrieval insertion. We conduct another user study with a third “Retrieval Appropriateness” question to gauge semantic appropriateness of the generated gesture during the insertion window (Fig. 8). Interestingly, we observe increasing preference of our retrieval-guided model as we move towards both extremes of the insertion. In terms of gesture perception, evaluators neither prefer generation without RAG (purely data-driven) nor a strong insertion through Inpainting (purely retrieval-based). Lastly, our model achieves good balance of retrieval following in terms of MPJPE, prosodic alignment and diversity.
@@
No RAG $0.519$ $0.447$ $8.64$ $59.6$
RAG by Inpainting [41] $0.446$ $0.450$ $8.46$ $28.1$
RAG by LI only $0.589$ $0.486$ $9.29$ $54.6$
RAG-Gesture (LI+RG) $0.447$ $0.471$ $9.04$ $35.0$